1. Angelo Jogged 10 times in 30 mins around the track with a radius of 50m. find his average speed and average velocity.


1
Expert's answer
2021-10-07T13:49:21-0400

"\text{Distance jogged=}10\times{2\pi\times{50}}=3142m\\
time=30mins=1800s\\
\text{Average speed=}\dfrac{Distance}{Time}=\dfrac{3142}{1800}=1.75ms^{-1}""Displacement=0m\\
Velocity=\dfrac{Displacement}{Time}=0ms^{-1}"